A 1.5 kg block of wood and a 2.5 kg block of steel are placed in thermal contact.
If the wood was at an initial temperature of 50°C and the steel was at an initial temperature of 15°C then what is the final equilibrium temperature of the wood and steel? (to 2 s.f and in °C)
[cwood = 2400 J kg−1 K−1, csteel = 490 J kg−1 K−1]
1
Expert's answer
2020-05-11T20:18:52-0400

The final equilibrium temperature is T.

The wood loses heat, the steel gains the heat. The amount of heat in both cases is the same according to equilibrium condition.

"Q(loss) = Cm(T_w - T) = 2400*1.5*(50-T) = 180000 - 3600T"

"Q(gain) = Cm(T-T_s) = 490*2.5*(T-15) = 1225T - 18375"

"Q(loss) = Q(gain)"

"180000-3600T = 1225T - 18375"

"4825T = 198375"

T = 41.11 oC -- the final equilibrium temperature of the wood and steel